A random sample is drawn from a population of known standard deviation 11.3. Construct a 90% confidence interval for the population mean based on the information given (not all of the information given need be used).
a. n = 36, x-=105.2, s = 11.2
b. n = 100, x-=105.2, s = 11.2
1
Expert's answer
2019-12-18T12:08:18-0500

a.

"90\%CI=(\bar x-1.645\frac{s}{\sqrt{n}},\;\bar x+1.645\frac{s}{\sqrt{n}})=\\
=(105.2-1.645\frac{11.2}{\sqrt{36}},\;105.2+1.645\frac{11.2}{\sqrt{36}})=\\
=(102.13,\;108.27)."


b.

"90\%CI=(\bar x-1.645\frac{s}{\sqrt{n}},\;\bar x+1.645\frac{s}{\sqrt{n}})=\\
=(105.2-1.645\frac{11.2}{\sqrt{100}},\;105.2+1.645\frac{11.2}{\sqrt{100}})=\\
=(103.36,\;107.04)."